### v1

# Topological and Graph Theoretical Analysis of Dynamic Functional Connectivity for Autism Spectrum Disorder

Abstract
Autism Spectrum Disorder (ASD) is a prevalent neurological disorder. However, the multi-faceted symptoms and large individual differences among ASD patients are hindering the diagnosis process, which largely relies on subject descriptions and lacks quantitative biomarkers. To remediate such problems, this paper explores the use of graph theory and topological data analysis (TDA) to study brain activity in ASD patients and normal controls. We employ the Mapper algorithm in TDA and the distance correlation graphical model (DCGM) in graph theory to create brain state networks, then innovatively adopt complex network metrics in Graph signal processing (GSP) and physical quantities to analyze brain activities over time. Our findings reveal statistical differences in network characteristics between ASD and control groups. Compared to normal subjects, brain state networks of ASD patients tend to have decreased modularity, higher von Neumann entropy, increased Betti-0 numbers, and decreased Betti-1 numbers. These findings attest to the biological traits of ASD, suggesting less transitioning in brain dynamics. These findings offer potential biomarkers for ASD diagnosis and deepen our understanding of its neural correlations.
Keywords:
Autism Dynamic functional connectivity Graph signal processing Topological data analysis
1 Introduction
Autism Spectrum Disorder (ASD) is a neurodevelopmental disability that has become more prevalent throughout the years. According to the study [12] of the Autism and Developmental Disabilities Monitoring (ADDM), in the year 2020, 1 in 36 children at the age of 8 was estimated to have ASD, higher than previous estimates during 2000-2018. Patients exhibit symptoms such as social interaction impairments and communication deficits, causing challenges for them to adapt to society. This makes the early diagnosis of this disease very important. However, the symptoms of ASD can vary in a wide range, including social, emotional, and cognitive impairments [8]. The lack of quantitative measures impedes the diagnosis process, calling for more advanced data analysis algorithms to uncover more accurate biomarkers of such diseases and bolster timely detection and intervention.
Brain network analysis tackles these problems. It is the study of the network relationship patterns for brain activities. By analyzing brain signals extracted from functional Magnetic Resonance Imaging (fMRI) or electroencephalography (EEG) data, these studies aim to uncover the spatiotemporal dynamics of human brains and facilitate disease analysis. Specifically, temporal information has gained great emphasis in recent studies. Compared with stationary patterns, dynamic signals can provide richer details of the brain [2, 11, 10]. Many disease patterns and natural phenomena can be uncovered by examining the time-varying brain signals. The study by Damaraju et al. [5] found that schizophrenia patients, on average, are less involved in large-scale brain activities, while they display more abnormal activities. However, one of the major challenges of studying brain activities is the immense amount of data involved, which awaits powerful algorithms to extract meaningful patterns and reduce the computation burden [3]. It is thus crucial to discern the underlying topology and structure of the data.
Topology is the study of shapes and structures. Topological data analysis (TDA) is an emerging field that aims to study the shape of data. A recently proposed tool in TDA is the Mapper approach [20], which reduces a set of high-dimensional data into a graph. In this process, the shape and topology of the original data samples are captured and represented by the graph. The Mapper algorithm is suitable for depicting the data distribution across subjects or time, and it has seen previous success. Saggar et al. [19] applied the Mapper algorithm to study brain activities at rest and under different tasks and found connectivity differences for different brain states. In their later work, they applied it to study the resting state fMRI [18] and uncovered important transition hub states using the node degree metrics. While they have made impactful discoveries, the analysis was limited to healthy subjects, and the exploration of graph metrics remained on the basic node degrees. It is thus interesting to adopt the Mapper algorithm in the context of brain diseases and explore different metrics on the Mapper graphs to uncover the biomarkers of brain alterations.
Graph signal processing (GSP) is an important technique in graph theory [16]. It extends the concepts in conventional digital signal processing to process data defined on graph domains. Basic GSP tools include sampling, filtering, and graph learning. GSP has a wide range of applications for brain network analysis. Huang et al. [9] decompose brain signals according to different smoothness (or rapidness) levels by using graph spectral operations and discovered the correspondence between brain network activities and graph frequencies. Preti et al. [17] leverage the structure-function coupling implication of graph frequency components, linking brain functional activities with neural architectures through node- and edge-level graph metrics. However, these methods construct graphs based on brain regions while the time information is collapsed in the correlation calculation step, hindering the detailed analysis of brain state distributions across time. Regarding this, the Mapper algorithm has clear advantages of not collapsing the temporal resolution. It is then reasonable to combine GSP with the Mapper algorithm to uncover the intricate brain dynamics across time.
This paper studies the brain activity distributions across time for autism and normal people. By employing the Mapper algorithm, a graph representation of the brain state distribution landscape can be created for each subject. Then, by studying the GSP and TDA properties of these graphs, the differences between the normal and autism groups are discerned. In summary, the contribution and novelty of this paper are:
  • To the best of the authors’ knowledge, it is the first work that applies the Mapper algorithm to study patients with neurological disorders.
  • This work inventively combines GSP with TDA, applying a variety of GSP metrics (including node-, graph-level, and frequency domain metrics) and TDA quantities to examine brain state distribution networks.
  • The group differences identified in the graph metrics reflect meaningful biological traits of altered brain activities induced by ASD. These findings attest to essential symptoms of ASD.
2 Methodology
Given the neuroimaging data, from which time-varying vectors can be extracted, we apply graph theory and TDA techniques to convert it into graph data for brain network analysis. Then, the networks are examined in terms of their structure, connectivity, and topology using various graph metrics. By comparing the metric results of normal and autistic groups, we study the differences in dynamic functional connectivity induced by autism.
2.1 Mapper algorithm
Figure 1: Diagram of the Mapper algorithm.
To reflect the brain state distribution at the individual level, we employ the Mapper algorithm from the field of Topological Data Analysis (TDA). The Mapper algorithm encapsulates the topology of data samples as a graph structure, providing an essential depiction of the data distribution landscape. This approach makes few assumptions about the data and is robust to noise due to its use of the graph structure.
As illustrated in Fig.1, the brain signal at each time point is viewed as a data vector in the data domain. The first step of the Mapper algorithm is to project data to a low-dimensional space where nearby points indicate similar brain activation patterns. We use principal component analysis (PCA) to achieve this step. After that, the dimension-reduced data space is divided into overlapping bins, and partial clustering is conducted inside each bin to combine similar data samples (i.e., time frames). In the final step, a graph is constructed by using the clusters as nodes, and an edge is established between two nodes if the corresponding clusters have one or more common data samples. Through this process, we obtained a brain graph dataset, which enabled us to explore its topological and physical characteristics.
In the realization of Geniesse et al. [7], which is adopted in our experiment, the graph is constructed as a reciprocal kNN graph, with an edge established between two nodes if they have common neighbors among the top-k nearest neighbors. This method stabilizes the noise level by reducing "false" connections.
Through the binning process in the data sample space, the Mapper graph can be viewed as a "histogram" of instantaneous brain states. Hence, it depicts the distribution (probability) of a subject’s brain activities across time. The Mapper algorithm encapsulates the topology of data samples as a graph structure, providing an essential depiction of the data distribution landscape. This approach makes few assumptions about the data and is robust to noise due to its use of the graph structure.
The rightmost figure in Fig. 1 shows an example of the graph created by the Mapper algorithm. One important quality of the Mapper algorithm is that through the clustering process, similar time points can be merged as the same node, thus capturing more complex organizations rather than a plain chronical chain, as illustrated in Fig. 2.
Figure 2: Formation of high-level interactions in the Mapper algorithm. Grouping similar signals in different time points enables the algorithm to model complex organizations in time.
2.2 Distance correlation graphical model
Distance correlation graphical models (DCGMs) are used in addition to the Mapper algorithm to model brain state networks. One of the most valued properties of the DCGM is that, because it does not have a PCA dimension reduction step, it treats every time point independently throughout the entire process, which can keep the chronological information intact and is vital to time-varying brain activity analysis.
As Fig. 3 shows, we construct the Distance correlation graphical model (DCGM) by calculating the distance correlation between each pair of time points. We then filter the correlation values using a p-value threshold, screening out the edges with p-values larger than 0.05 to ensure that only statistically important connections are kept in the graph. The resulting graph has each node as a time point, and each edge represents the correlation between the time points on its two ends.
Figure 3: Diagram of the distance correlation graphical model.
Distance correlation measures the association between random variables by comparing the pairwise distances in the joint distribution of variables to their marginal distributions. For two random variables $X$ and $Y$, the distance correlation between them is defined as follows:
$\text{dCor}(X,Y)=\sqrt{\frac{{\text{dCov}^{2}(X,Y)}}{{\text{dVar}(X)\text{dVar}(Y)}}}$ (1)
where $dCov$ is the distance covariance, and $dVar$ is the square root of the product of distance variances.
DCGMs capture nonlinear relationships between variables, which is a limitation of traditional linear methods like covariance or partial correlation. This allows DCGMs to model complex dependencies more accurately. Distance correlation is less sensitive to outliers compared to Pearson correlation, making DCGMs more robust in the presence of noisy data.
2.3 Modularity and graph eigenvalue
We introduce the graph modularity metric to examine the structure of the individual graphs generated by the Mapper algorithm. We also perform frequency domain analysis on the eigenvalue distributions of the graphs. These metrics are introduced below:
Modularity: Modularity assesses the presence of community structures within the network, identifying groups of nodes that exhibit higher connectivity with each other than with the rest of the network. This metric measures the degree to which a network can be divided into modules, which unveils the modular organization of the brain’s temporal state arrangement. Its mathematical expression is:
$Q=\frac{1}{2m}\sum_{ij}\left(A_{ij}-\frac{k_{i}k_{j}}{2m}\right)\delta(c_{i},c_{j})$ (2)
where $A_{ij}$ are elements of the adjacency matrix, $k_{i}$ is the Degree of node $i$, $m$ is the total number of edges in the graph, $c_{i}$ and $c_{j}$ are Community assignments of nodes $i$ and $j$, and $\delta(c_{i},c_{j})$ is the Kronecker delta, which is equal to 1 if $c_{i}=c_{j}$ and 0 otherwise.
Laplacian Eigenvalues: Eigenvalue distributions of the graph’s Laplacian matrix offer insights into the overall network dynamics. Features such as the mean or variance of eigenvalue provide information about stability and the presence of distinct functional modules. We can obtain a Laplacian matrix $L$ by $L=D-W$, and the degree matrix $D\in\mathbb{R}^{n\times n}$ is a diagonal matrix with diagonal elements where $D[i,i]=\sum\limits_{j\neq i}W[i,j]$ represents the weighted degree of the nodes. It can be proved that $L$ is real, symmetric, and positive semi-definite. After obtaining the Graph Laplacian $L$, we can decompose it using eigendecomposition $L=U\Lambda U^{H}$ such that the diagonal matrix $\Lambda$ has entries:
$\lambda_{i}=\{0=\lambda_{0}\leqslant\lambda_{1}\leqslant\ldots\leqslant\lambda_{n-1}\}$ (3)
which will be the Laplacian eigenvalues of the network. Generally, lower eigenvalues correspond to smoother signals across the graph, where there is less variation among node values, while higher eigenvalues indicate the signal components that change abruptly among nodes [16].
Von Neumann entropy: Building on the definition of Laplacian eigenvalues, we can characterize the degree of randomness of a graph structure as the von Neumann entropy of its Laplacian eigenvalues. The mathematical formulation is then:
$S=-\sum_{\lambda_{i}\neq 0}{\frac{\lambda_{i}}{len(\lambda)}\times\log{\frac{\lambda_{i}}{len(\lambda)}}}$ (4)
where $\lambda$ is the set of all eigenvalues and $\lambda_{i}$ is the $i^{th}$ eigenvalue
2.4 Topological analysis
To examine the brain state distributions of the subjects, we explore the structural aspect of the simplicial complex composed of instantaneous brain state data points.
Betti number: In topological analysis, a Vietoris-Rips complex can be constructed by selecting a distance threshold, building edges between data points that are less distant than the threshold, and creating higher-level topological structures wherever possible. Then, the numbers of high-dimensional shapes are counted as the characteristics of the simplicial complex, which is annotated as the Betti numbers.
2.5 Ollivier-Ricci curvature
Proposed by Ollivier [15], the Olliview-Ricci curvature describes the connectivity and topology of each node in the graph. For node $x$, $y$ $\in V$ in G(V,E), their Ollivier-Ricci curvature can be defined as following:
$ORC(x,y)=1-\frac{W\left(m_{x},m_{y}\right)}{d(x,y)}$ (5)
where $W\left(m_{x},m_{y}\right)$ is the $L^{1}$ Wasserstein distance and $d(x,y)$ is the shortest path distance between $x$ and $y$. According to Chien-Chun Ni [14], we can calculate $m_{x}$ and $m_{y}$ as measures of mass distributions of $x$ and $y$ by lazy random walks on graphs.
3 Experiment and result
3.1 Data preparation
The experiment data includes the study of brain state distribution differences between autistic and normal people. By using the Nilearn package [1] in Python, the pre-processed data in the ABIDE dataset [4] can be obtained. The pre-processed steps include bandpass filtering, global signal regression, and quality check, ensuring a higher data quality. The fMRI data is then registered and masked using the icbm152 2009 [6] to extract time series. These pre-processing steps are crucial in brain data analysis, as factors like individual differences and motion noise can lead to data ambiguity. As is shown in Fig. 4, the pre-processing and quality of data can make a profound impact on the analysis in the later stage.
Figure 4: A comparison of graphs created with pre-processed and unprocessed data. (A). Graph created using data with bandpass filtering or global signal regression. (B). Without bandpass filtering or global signal regression. The parameters for graph creation in both subfigures are the same, as mentioned in Section 3.2.
3.2 Network construction
To create the networks that reflect the brain state distribution at the individual level, we examine both the Mapper algorithm and the Distance correlation graphical model.
Fig. 4 (A) shows an example of the graph created by the Mapper algorithm. Some parameters in the Mapper algorithm can affect the resulting graph . After taking note of the parameter searching process of Saggar et al. [19] and testing out different choices, we set the parameters as below in this experiment:
  • Resolution ($R$): 50
  • Gain ($G$): 30%
  • Number of neighbors ($k$): 6
  • Dimension reduction method: PCA with 50 PCs
For the construction of the DCGM, we first calculate the distance correlation between brain signal vectors at different time points. Then, we screen out the statistically insignificant ones using a p-value threshold of 0.05, thus obtaining the weighted adjacency matrix, as shown in Fig. 3.
3.3 Graphical and topological metrics plots
As we examine the traits and patterns of the brain-state graphs generated by the Mapper algorithm and the DCGM, the findings and comparisons are plotted in Fig.5. The violin plots illustrate the distributions of metric values in autistic and control groups, with box plots included inside the violin plots indicating the median and quartiles of the distributions. The histograms show the detailed distributional features of the two groups’ metric values and their side-by-side comparisons. By observing these plots, including the peaks, tails, quartiles, and shapes of these distributions, we can make the following analysis of the data:
Figure 5: Results of graph and topological metrics for autism and normal groups. The left column (a, c, e, g, i) shows the violin plots and histograms of modularity, average eigenvalue, entropy, average curvatures, and average curls of DCGM graphs, while the right column (b, d, f, h, j) shows these metric results of the Mapper graphs.
Modularity: As shown in Fig. 5(a-b), in graphs created by either method, there is a tendency for decreased modularity in the autism group. These plots show that, compared with normal controls, autism patients tend to have lower network modularities on average, indicating less clear-cut community structures in the network. This contrast implies that ASD patients display less distinct types of brain activities or fewer transitions in brain states at rest. This result reflects that the autism subjects have less distinct "community" or "clusters" of brain states, which can indicate the overly smooth brain activity discovered in [21].
Average eigenvalue: The frequency domain analysis is conducted on both Mapper and DCGM graphs by studying the eigenvalue distribution, as is indicated in Fig. 5(c-d). Interestingly, the average eigenvalues of DCGM and Mapper graphs lead to contradicting results. But overall, there are no significant differences between the average eigenvalues of normal and autistic groups.
According to the theory of GSP [16], eigenvectors of smaller eigenvalues correspond to smoother graph signals across nodes. However, comparing eigenvalues themselves may not have direct implications. Considering the temporal graph’s ability to characterize brain state distributions across time, it can be valuable to examine the frequency components of temporal graph signals, as it can reveal brain dynamics over time.
Von Neumann entropy: The networks of autism subjects are likely to have higher von Neumann entropy. The DCGM captures this phenomenon more clearly. The increased entropy values imply more random states of mind in autistic subjects at rest. Together with the comparison of network modularity, this shows that the autistic subjects have less organized, distinct brain activities at rest, which could be an indicator of restricted interests and repetitive behaviors as introduced in [8].
Node curvatures: Fig. 5(g-h) indicates the distributions of node curvature values in normal and autistic groups. Although the difference between autism and normal groups is not remarkable, the distribution of curvature scores for the two graph creation methods is very distinct. These plots show that compared to DCGM graphs, Mapper graphs have much more variant node curvature values. This implicates that the Mapper algorithm can create relatively more diverse graph structures.
Simplicial weights curl: Simplicial weights curl depicts the flow of weights of the simplicial complex constructed from the brain network. According to Fig. 5(i-j), the Mapper graph and the DCGM yield different curl comparisons for normal and autism groups. One way to make sense of this is that the simplicial weights curl is related to the methods of network creation. Hence, future work will consider simplicial weights curl as structural criteria to guide the network generation process and select the optimal network generation method that best reflects subjects’ time-varying brain activity.
Topological Analysis In our experiment, the topological analysis is realized by the GUDHI package [13] in Python. The simplicial complex is created using a maximum edge length of 500. We collect the zeroth- and first-order Betti numbers of these patients, which are shown in Fig. 6. As can be observed in the plots, compared to normal people, autistic subjects tend to have higher counts of $\beta_{0}$ numbers and decreased counts of $\beta_{1}$ numbers, indicating a larger number of scattered components and a lower amount of complex interaction structures.
Figure 6: Betti numbers $\beta_{0}$ and $\beta_{1}$ for brain states of autism and normal subjects under the same threshold (distance $\leq$ 550).
4 Conclusions and future work
This study used topological data analysis (TDA) and compared the complex network metrics to investigate the difference in brain activities between individuals with autism spectrum disorder (ASD) and neurotypical individuals. By applying the Mapper algorithm and the distance correlation graphical model (DCGM), we constructed brain state distribution networks for each subject, enabling a deeper understanding of the underlying patterns in brain complex networks.
Both algorithms are accompanied by graph theory and topological data analysis to investigate the temporal pattern of brain connectivity. The application of temporal graphs allowed us to have a deeper understanding of the human brain’s temporal transitions under various conditions. The results of graphical and topological metrics implicated less distinct and more random brain states for autistic subjects, corroborating previous findings in neuroscience studies and proving the ability of graph and topological analysis to capture the alteration in brain activity of neurological disorders.
To take further steps, we plan to substantiate this discovery on a more extensive dataset of autism spectrum disorder. The devised method can also be applied to other neuroscience phenomena, especially the ones that are known to have alterations in brain state temporal sequences. By studying other neurological events, we seek to provide insights into the underlying commonality and differences among these neuroscience conditions. To design a more precise and detailed analysis, we plan to narrow down the analysis to the specific brain regions that have been proven to cause the neurological condition in the study, which we expect will enhance the contrast between groups.
Lastly, considering the extensive use of matrix eigenvalues in signal processing, statistical physics, and complex network analysis for analysis and simulation of network evolution, we recognize the potential of transplanting and applying these techniques to brain complex networks. These physical quantities may offer better descriptions of the developmental processes of psychiatric disorders. Particularly for non-stationary processes like brain signals, more time-series algorithms need to be incorporated to analyze such complex network evolution.